A lateral edge of the hexagonal pyramid is 20m long and its base perimeter is 6m determine the volume of the pyramid and its lateral area
1
Expert's answer
2020-03-28T14:29:34-0400

V = 1/3 * S * H

S = 1/2 * a 2  * 3 *"\sqrt {3}"

a = P / 6 = 6/6 = 1 m

S = 1.5 * "\sqrt {3}" m2

H = "\sqrt{400-1}" = "\sqrt{399}" m

V = 1/3 * 1.5 * "\sqrt {3}" * "\sqrt{399}" = 1.5 "\sqrt{133}" m3

h = "\sqrt{400-0.25}" = 0.5 * "\sqrt{1599}" m

 = 0.5 * P * h = 1.5 * "\sqrt{1599}" m2

The answer to the problem: V = 1.5 "\sqrt{133}" m3 , Sl  = 1.5 * "\sqrt{1599}" m2